# A Review of Graph Neural Networks in Epidemic Modeling

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXX
@@ 3.1.2. Surveillance
To illustrate, tasks like detecting infected individuals promptly (songCOVID19InfectionInference2023) and estimating infection risks in different locations in real-time (wang2023wdcip; gouarebDetectionPatientsRisk2023; hanDevilLandscapesInferring2023) can be seen as surveillance tasks, as they their prediction targets lie in present or near future.
@@ 4.1.2. Temporal Dynamics Modeling
where h is simply the output of an MLP (Multilayer Perceptron) over the node temporal features x at time $t$ reaching back $d$ days. Another surge of approaches first executes graph spatial convolution in each time step separately, and then leverages all outputs of the GNNs as the input of the temporal module and utilizes them for final downstream tasks like the prediction (gao2021stan; sijirani2023spatiotemporala; moon2023graph; duarte2023time). STEP (sijirani2023spatiotemporala) execute the multi-layers graph convolution operation to get all node embedding h, and then leverage the GRU to get the final output:

### v4

@@ Abstract.
copyright: acmlicensedjournalyear: 2024copyright: acmlicensedconference: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’24), August 25–29, 2024, Barcelona, Spaindoi: 10.1145/3637528.3671455isbn: 979-8-4007-0490-1/24/08
@@ 2.3. Mechanistic Models
Furthermore, utilizing the SIR model within the context of graph-based structures leads to the development of what is referred to as the Network SIR model (sha2021source; brede2012networks; balcan2009multiscale; venkatramanan2017spatiotemporal). Within this framework, an infectious node $j$ has the potential to transmit the infection to another node $i$, provided that $i$ is susceptible and located adjacently to $j$ (i.e., $i\in A(j)$ ). The probabilities of the node $i$ being in a susceptible $S_{i}(t)$, infected $I_{i}(t)$, or recovered $R_{i}(t)$ state at any given time $t$, they are recalculated as follows:
(4) $\begin{gathered}\frac{dS_{i}(t)}{dt}=-\beta\sum_{j}A_{ij}S_{i}(t)I_{j}(t),\\ \frac{dI_{i}(t)}{dt}=\beta\sum_{j}A_{ij}S_{i}(t)I_{j}(t)-\gamma I_{i}(t),\quad\frac{dR_{i}(t)}{dt}=\gamma I_{i}(t).\end{gathered}$
Considering the initial phase where $S_{i}(t)$ is nearly one, the model facilitates the calculation of the basic reproduction number $R_{0}$ as $R_{0}=\beta\lambda_{1}/\gamma$, with $\lambda_{1}$ representing the leading eigenvalue of the adjacency matrix $A$.
@@
Bridging the Gap. To bridge the gaps between mechanistic models and neural models like GNNs, a Python library called EpiLearn  (liu2024epilearn) is also developed at https://github.com/Emory-Melody/EpiLearn. This library serves as a toolkit for epidemic modeling and analysis, empowering data mining on epidemiology data with machine learning models.
@@
GraphDNA (yang2022dynamic) GCN + LSTM Spatial-Temporal Graph; Dynamic Graph Structure
@@ 3.1.2. Surveillance
To illustrate, tasks like detecting infected individuals promptly (songCOVID19InfectionInference2023) and estimating infection risks in different locations in real-time (wang2023wdcip; yang2022dynamic; gouarebDetectionPatientsRisk2023; hanDevilLandscapesInferring2023) can be seen as surveillance tasks, as they their prediction targets lie in present or near future.
@@ 4.1.2. Temporal Dynamics Modeling
where h is simply the output of an MLP (Multilayer Perceptron) over the node temporal features x at time $t$ reaching back $d$ days. Another surge of approaches first executes graph spatial convolution in each time step separately, and then leverages all outputs of the GNNs as the input of the temporal module and utilizes them for final downstream tasks like the prediction (gao2021stan; yuSpatiotemporalGraphLearning2023; moon2023graph; duarte2023time). STEP (yuSpatiotemporalGraphLearning2023) execute the multi-layers graph convolution operation to get all node embedding h, and then leverage the GRU to get the final output: